Session: Nov 1998 (session 4)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (M. Bondi at Fri Nov 27 02:56:45 MET 1998) Success

Shanghai (Huang Xinyong at Fri Nov 27 06:32:44 MET 1998) Success
The exp. has finished. The system seems OK. But the LO of the BBC05 is not good. The system noise is not correct in the start of the log file. Tape was sent out from Shanghai to Bonn by DHL on 27 nov.

Torun (K.Borkowski at Fri Nov 27 09:22:37 MET 1998) Success
See however a note for the session.

Jodrell (Alastair Gunn at Fri Nov 27 12:04:52 MET 1998) Success
Strong RFI. No known problems.

Cambridge (Alastair Gunn at Fri Nov 27 12:05:29 MET 1998) Failure
Experiment cancelled for Cambridge 32-m telescope due to serious hardware fault with VLBA recorder head stack.

Effelsberg (dgraham at Fri Nov 27 13:17:55 MET 1998) Success
seems OK

Onsala (John Conway at Sat Nov 28 17:59:07 MET 1998) Success
No known problems

Noto (C.Trigilio at Mon Nov 30 11:14:18 MET 1998) Success

Effelsberg (dgraham at Mon Nov 30 15:12:23 MET 1998) Minor failures
ignore previous report ! there was a break for snowfall 1830-2000 !

Westerbork (Tony FOley at Tue Dec 1 18:23:33 MET 1998) Minor failures
Scan before 19:30UT are outside Hour Angle range

Calibration (Jean\-Francois Desmurs at Fri Dec 11 16:33:34 MET 1998) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the nov98 directory under the name: c98l4.antab. Read the comments in the file to get further information.